# Stablecoin Runs and Disclosure Policy

### v1

@@ 4.1 Case Studies
  • Circle and the Collapse of Silicon Valley Bank. Since USDC is a fiat-backed stablecoin whose issuer, Circle, makes frequent public disclosures about reserves which are generally trusted, USDC would correspond to a stablecoin with a low prior variance. Our model predicts good financial stability when fundamentals are good and poor financial stability when fundamentals are poor, relative to coins of a similar reserve structure but different levels of transparency. During SVB’s collapse, when Circle disclosed that SVB held a portion of their reserves, USDC depegged shortly after. Upon the announcement by the Federal Deposit Insurance Corporation (FDIC) that SVB would be bailed out, USDC repegged not long after (see Figure 5). In this case, as a consequency of transparency, collateral risk is a major determinant of financial stability, as opposed to large sale risk.
@@
  • Tether’s Disclosures and Depegging Events. For Tether, disclosures of reserves are less frequent and less verifiable due to operating outside of a regulatory framework, resulting in a higher prior variance. Compared to more transparent stablecoins, there is more exposure to run risk when fundamentals are good, but less exposure to large sale risk when fundamentals are lower but not too low (see the intermediate interval in Figure 2). Additionally, the asymmetric information generated by Tether’s transfers and recollateralizations corresponds to a large dispersion in private signals, which may additionally reduce run risk when fundamentals are good. These reasons could explain why despite the opaqueness and occasional depegging driven by large sales, Tether repegs in a short amount of time, avoiding more substantial depegging incidents or a full collapse.
@@
  • Terra–Luna Collapse. Endogenously-backed stablecoins like UST tend to have high prior variances due to their self-referential nature. For example, Ahmed, Aldasoro, and Duley (2024) find empirical evidence that Luna, the cryptoasset backing TerraUSD, had a non-constant conditional volatility, leading to unconditional fat tails. Additionally, our model suggests that single-collateralized stablecoins, including UST, should generate less heterogeneous beliefs and thus more precise private signals, heightening large sale risk when fundamentals are good. As the collapse of UST and Luna started with a large sale, possibly of speculative nature, it could be that a combination of relatively large prior variance and precise private signals contributed to the full collapse.
Our discussion adds on to that of Ahmed, Aldasoro, and Duley (2024). They discuss the effect of public information and fat-tailed distributions on real-world depegging events, which we complement by (i) attributing the causes of events to either poor collateral or large sales and (ii) considering the affect of the variation in private signals on the events. In line to the above analysis, we also offer a rough characterization of the common prior variance and private signal variance based on the type of stablecoin.
Stablecoin Type Common Prior Variance $(\sigma)$ Private Signal Variance $(\sigma_{x})$
Transparent Fiat-Backed Low Low–Medium
Opaque Fiat-Backed High Medium–High
Single-Collateral Crypto-Backed Varies Low–Medium
Multi-Collateral Crypto-Backed Varies Medium–High
Algorithmic High Varies
Table 3: Classification of common prior and private signal dispersion based on stablecoin type.
Note that for real-world stablecoin designs, a large common prior variance is correlated with a large private signal variance. In the case of opaque fiat-backed stablecoins, the variance of the common prior arises from the issuer’s opaque disclosure policy, while the variance of the private signal arises from the information asymmetry generated from said policy. Also worthy of note is how the basket of collateral reserves affects parameters for crypto-collateralized stablecoins. Baskets consisting of anti-correlated cryptoassets have less systemic risk, and thus a less common prior variance, while the opposite is true for single-collateralized tokens or baskets with correlated assets. As investors may have different beliefs about each asset’s returns, having a basket of diverse assets may also widen the heterogeneity among investors, thus raising the private signal variance.
4.2 Possible Benefits of Opaqueness
The model results, case studies, and table above suggest that in certain contexts, there may be a benefit to opaqueness, whether that comes in the form of an increase in $\sigma$ (common prior) or $\sigma_{x}$ (private signal).
With regards to the dispersion of the common prior, Proposition 3.2 and Figure 2 suggest that there is an interval, in which fundamentals are neither too good or poor, where large sale risk decreases and collateral risk increases in variance. Based on the large sale probability and difference between the critical thresholds ($\bar{\theta}_{1}(x^{\star}_{1})$ and $\bar{\theta}_{0}(x^{\star}_{0})$), the overall run probability may decrease in variance as well. For instance, when there is a moderate collateral shock to drive $\mu$ in that interval, higher variance may shield the issuer from large sales, especially those of speculative nature. Nondisclosure or avoidance of traditional financial institutions for deposits may also shield stablecoins from shocks to those institutions. For example, Tether remained relatively stable during the SVB crisis by not having any exposure to SVB. This, coupled with a less transparent disclosure policy, suggests that a risk-dominant equilibrium where investors gravitate towards holding arises.
With respect to the dispersion of the private signal, Proposition 3.3 and Figure 3 suggest that a larger dispersion is beneficial for financial stability when fundamentals are good. To this end, a larger user base of the stablecoin or a more diverse basket of collateral reserves may achieve a more heterogeneous user base. Ahmed, Aldasoro, and Duley (2024) run an event study when Dai, a leading crypto-collateralized stablecoin, switched from a single-collateral to a multi-collateral structure. They find that Dai’s peg stability improved, and explain this as a reduction in the volatility of the common prior as a result of risk diversification. However, risk is not necessarily diversified when the collateral assets are correlated, as may be the case for cryptocurrency. Our extended analysis posits that more contrasting investor beliefs could be a factor towards Dai’s improved peg stability.
The main policy implication for regulators is that while enforcing regulations on disclosures may improve trust in the DeFi ecosystem, disclosures would make shocks to the TradFi ecosystem spill over more easily into the world of DeFi, especially given the prominence of fiat-backed stablecoins whose issuers hold bank deposits. To build a resilient overall financial ecosystem, regulators must carefully consider and balance these two incentives. For stablecoin isusers, the main design implication is that the selling pressure of large sales and the diversification of collateral reserves should be considered when choosing what reserves to hold and how to go about disclosing them.
4.3 Testable Implications
The discussion so far lends itself to two empirically testable implications.
  • The Effect of Large Sales on Peg Stability. Proposition 2.1 suggests that a large sale increases selling pressure on investors. Empirically, we should thus observe an adverse effect on peg stability immediately following a large sale for a short amount of time. As large sales usually follow collateral shocks, it would be interesting to conduct the analysis in times where fundamentals are good.
  • The Effect of Collateral Reserves on Peg Stability for Crypto-Collateralized Stablecoins. While the number of fiat-backed stablecoin issuers who disclose reserves is likely not large enough for a thorough cross-sectional analysis, crypto-collateralized stablecoins have their reserves locked in on-chain smart contracts. We can then regress peg stability on weighted reserve volatility and a metric measuring reserve diversification. Our model predicts that peg stability should decrease in reserve volatility and increase in reserve diversification.
4.4 Distance to No-Questions-Asked
We conclude by discussing the differences between stablecoins and money using the no-questions-asked framework of Holmström (2015), describing money as a information insensitive and deep-in-the-money debt contract, i.e. something which can always be redeemed at par. While stablecoin prices behave more similarly to pegged currencies than bonds, there is a link between our model and the notions of information insensitivity and deep-in-the-moneyness. Reserve assets that are “deep-in-the-money” could correspond to a prior distribution that makes it unlikely for $\theta$ to be realized below the critical thresholds (e.g. high $\mu$, relatively low $\sigma$).
Holmström (2015) measures information sensitivity using a metric called information acquisition sensitivity (IAS), roughly described as the benefit that acquiring information has on making a decision (in our case, to sell or to hold). A low common prior dispersion implies a low IAS since as signals become more concentrated around the true value, the relative benefit of information is reduced. A high private signal dispersion also implies a low IAS due to the risk-dominance of the equilibrium strategy: as other investors are more likely to choose the “safe” option, knowing more information about the true value of reserves is less likely to change the equilibrium outcome. The attributes of parameters in our model that are desirable for financial stability are thus linked to desirable properties of money as described in Holmström (2015).
@@ 5 Conclusion
Appendix A Proofs
A.1 Proof of Proposition 2.1
Suppose that there is no large sale. Observe that $\bar{\theta}_{0}$ is the unique value of $\theta$ solving
$\theta-(1-\delta)(1-F_{\bar{x},\sigma_{x}}(\theta))=0$
and $x^{\star}_{0}$ is any value of $\bar{x}$ that solves
$F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))-\frac{\eta}{1+\eta}=0.$
We first show that $F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))$ is decreasing in $\bar{x}$. By the Implicit Function Theorem,
$\frac{d\bar{\theta}_{0}}{d\bar{x}}=-\frac{(1-\delta)\cdot\frac{\partial}{\partial\bar{x}}F_{\bar{x},\sigma_{x}}(\bar{\theta}_{0})}{1+(1-\delta)\cdot\frac{\partial}{\partial\theta}F_{\bar{x},\sigma_{x}}(\bar{\theta}_{0})}=\frac{\frac{1-\delta}{\sigma_{x}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\bar{x}}{\sigma_{x}}\right)}{1+\frac{1-\delta}{\sigma_{x}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\bar{x}}{\sigma_{x}}\right)}>0.$
By the Chain Rule,
$\frac{d}{d\bar{x}}F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))$ $=\frac{\partial}{\partial\bar{x}}F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))+\frac{\partial}{\partial\theta}F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))\cdot\frac{d\bar{\theta}_{0}}{d\bar{x}}$
$=-\frac{\mu^{\prime}_{p}(\bar{x})}{\sigma_{p}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\mu_{p}(\bar{x})}{\sigma_{p}}\right)+\frac{1}{\sigma_{p}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\mu_{p}(\bar{x})}{\sigma_{p}}\right)\cdot\frac{d\bar{\theta}_{0}}{d\bar{x}}$
$=\frac{1}{\sigma_{p}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\mu_{p}(\bar{x})}{\sigma_{p}}\right)\cdot\left(\frac{d\bar{\theta}_{0}}{d\bar{x}}-\mu_{p}^{\prime}(\bar{x})\right),$
so its sign depends on $d\bar{\theta}_{0}/d\bar{x}-\mu_{p}^{\prime}(\bar{x})$. Observe that
$\frac{d\bar{\theta}_{0}}{d\bar{x}}-\mu_{p}^{\prime}(\bar{x})$ $=\frac{\frac{1-\delta}{\sigma_{x}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\bar{x}}{\sigma_{x}}\right)}{1+\frac{1-\delta}{\sigma_{x}}\cdot\phi\left(\frac{\bar{\theta}_{0}-\bar{x}}{\sigma_{x}}\right)}-\frac{\sigma^{2}}{\sigma^{2}+\sigma_{x}^{2}}$
$\leq\frac{\frac{1-\delta}{\sqrt{2\pi}\sigma_{x}}}{1+\frac{1-\delta}{\sqrt{2\pi}\sigma_{x}}}-\frac{\sigma^{2}}{\sigma^{2}+\sigma_{x}^{2}},$
since the maximum value of a standard normal density is $1/\sqrt{2\pi}$. It follows that if $(1-\delta)\sigma_{x}<\sqrt{2\pi}\sigma^{2}$, then $d\bar{\theta}_{0}/d\bar{x}-\mu_{p}^{\prime}(\bar{x})<0$. This holds by Assumption 1, so $F_{\bar{x},\sigma_{x}}(\theta)$ is decreasing in $\bar{x}$.
We check the limiting behavior of $F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))$. Note that $1-F_{\bar{x},\sigma_{x}}(\theta)\to 0$ as $\bar{x}\to-\infty$, so $\bar{\theta}_{0}\to 0$ as $\bar{x}\to-\infty$. Since $\mu_{p}(\bar{x})\to-\infty$ as $\bar{x}\to-\infty$ also, we have $F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))$ as $\bar{x}\to-\infty$. Conversely, note that $1-F_{\bar{x},\sigma_{x}}(\theta)\to 1$ as $\bar{x}\to\infty$, so $\bar{\theta}_{0}\to 1-\delta$ as $\bar{x}\to\infty$. Since $\mu_{p}(\bar{x})\to\infty$ as $\bar{x}\to\infty$ also, we have $F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{(}\theta_{0}(\bar{x}))\to 0$ as $\bar{x}\to\infty$. Since $F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{0}(\bar{x}))$ is decreasing (and continuous) with limits $1$ and 00 as $\bar{x}$ tends to $\infty$- ∞ and $\infty$, respectively, $x^{\star}_{0}$ exists and is unique.
Suppose that a large sale occurs. Recall that $\bar{\theta}_{1}$ is the unique value of $\theta$ solving
$\theta-\delta-(1-\delta)(1-F_{\bar{x},\sigma_{x}}(\theta))=0$
and $x^{\star}_{1}$ is any value of $\bar{x}$ that solves
$F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{1}(\bar{x}))-\frac{\eta}{1+\eta}=0.$
We first show that $F_{\mu_{p}(\bar{x}),\sigma_{p}}(\bar{\theta}_{1}(\bar{x}))$ is decreasing in $\bar{x}$. By the Implicit Function Theorem,
[…]

### v2

@@ 4.1 Case Studies
  • Circle and the Silicon Valley Bank Collapse. Since USDC is a fiat-backed stablecoin whose issuer, Circle, makes frequent public disclosures about reserves which are generally trusted, USDC would correspond to a stablecoin with a low prior variance. Our model predicts good financial stability when fundamentals are good and poor financial stability when fundamentals are poor, relative to coins of a similar reserve structure but different levels of transparency. During SVB’s collapse, when Circle disclosed that SVB held a portion of their reserves, USDC depegged shortly after. Upon the announcement by the Federal Deposit Insurance Corporation (FDIC) that SVB would be bailed out, USDC repegged not long after. In this case, as a consequency of transparency, collateral risk is a major determinant of financial stability, as opposed to large sale risk.
@@
  • Tether’s Disclosures and USDT Depegging Events. For Tether, disclosures of reserves are less frequent and less verifiable due to operating outside of a regulatory framework, resulting in a higher prior variance. Compared to more transparent stablecoins, there is more exposure to run risk when fundamentals are good, but less exposure to large sale risk when fundamentals are lower but not too low (see the intermediate interval in Figure 2). Additionally, the asymmetric information generated by Tether’s transfers and recollateralizations corresponds to a large dispersion in private signals, which may additionally reduce run risk when fundamentals are good. These reasons could explain why despite the opaqueness and occasional depegging driven by large sales, Tether repegs in a short amount of time, avoiding more substantial depegging incidents or a full collapse.
@@
  • Terra–Luna Collapse. Endogenously-backed stablecoins like UST tend to have high prior variances due to their self-referential nature. For example, [1] find empirical evidence that Luna, the cryptoasset backing TerraUSD, had a non-constant conditional volatility, leading to unconditional fat tails. Additionally, our model suggests that single-collateralized stablecoins, including UST, should generate less heterogeneous beliefs and thus more precise private signals, heightening large sale risk when fundamentals are good. As the collapse of UST and Luna started with a large sale, possibly of speculative nature, it could be that a combination of relatively large prior variance and precise private signals contributed to the full collapse.
  • GENIUS Act. The GENIUS Act’s requirements of full backing and transparent reporting significantly changes the environment in which US-based stablecoin issuers and investors interact. In our model’s terms, the act requires (i) stablecoin fundamentals to be consistently good and (ii) public information to be very precise. This act resolves the issues with Circle, which held risky cash deposits, by requiring high quality reserves, as well as the issues with Tether, which raised doubts with its dubious disclosure policies.
4.2 Information Design and Regulation
In our framework, greater dispersion in private signals can be a stabilizing force when reserve fundamentals are sufficiently strong, i.e. stablecoin issuers can benefit from a certain degree of heterogeneity in how investors interpret information. In practice, this does not mean obfuscation or misinformation, but rather design choices that naturally generate diverse views: cultivating a broad and heterogeneous user base (retail, DeFi protocols, market-makers, institutions), holding a diversified but fundamentally safe reserve portfolio, and communicating rich but non‐coarse disclosures that different investors process with different models.
The GENIUS Act, however, pushes in the opposite direction along several margins: by mandating conservative, simple reserve structures and frequent disclosures of reserve composition and redemption policies, it sharply increases the precision of public information, thus reducing room for informational heterogeneity that comes purely from opacity or selective disclosure. In a post-GENIUS environment, an issuer cannot “engineer” private signal dispersion by withholding or shading information; any attempt to do so would run into compliance and enforcement risk. Thus, US-based stablecoin issuers could be less capital efficient than offshore issuers not subject to the same regulations, who might hold riskier reserves or have opaque disclosures, but can effectively create stability via information design.
Our model implies that GENIUS tilts the information structure toward a regime where public information dominate and runs, when they occur, are more tightly tied to genuine deterioration in fundamentals. Residual dispersion in private signals coming from heterogeneous agents can still provide a buffer against self-fulfilling runs in good states, but is now a second-order issue. This highlights an important tension in the GENIUS framework: while it improves auditability and aligns runs more closely with fundamentals, it may do so at the cost of capital efficiency by mandating exclusively dollar-equivalent assets.
A more careful regulatory approach could consider alternative frameworks that preserve ex post auditability without making full public disclosure the default, such as continuous supervisory audit rights or cryptographic proof-of-reserves mechanisms that verify solvency without revealing custodial counterparties. While such approaches are likely second-order relative to the binding constraint on reserve composition, they illustrate how information policy, not just asset restrictions, shapes the competitive landscape between regulated and offshore stablecoin issuers.